OpenStudy (anonymous):

What does it mean that Z is the centroid and what does it tell us about segments (AN), (BO), and (CM)?

OpenStudy (anonymous):

where is the picture?

OpenStudy (anonymous):

here's the attachment

OpenStudy (anonymous):

Please get back to me when you get this.

OpenStudy (anonymous):

Centroid z is the point of concurrent of the three median lines. Median lines divides the opposite side of the triangle at its midpoint

OpenStudy (anonymous):

Base on the centroid theorem you will get AZ = 2/3 AN, BZ = 2/3 of B0, CZ= 2/3 CM

OpenStudy (anonymous):